### v1

# Robust Nonlinear Least Squares Problems

Abstract
This paper considers min-max optimization for a class of robust nonlinear least squares problems. We show via an example that solving the first order optimality conditions defined by gradients of the objective function can lead to incorrect solutions of min-max problems. We give an explicit formula for the value function of the inner maximization problem. Using the formula, we show that finding a first order $\epsilon$-approximate necessary minimax point of the min-max problem needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of the function value and gradients of the objective function. Moreover, we establish error bounds from any solution of the nonlinear least squares problem to the solution set of the robust nonlinear least squares problem.
keywords:
Min-max optimization, nonlinear least squares problems, robust learning, nonsmooth nonconvex optimization, complexity analysis
{AMS}
49K35, 90C30, 90C31
1 Introduction
[…]
$\min_{\mathbf{x}}\|F(\mathbf{x})\|^{2}$ (1)
is widely used in data analysis and machine learning. For example, Recurrent Neural Netwoks (RNN)[2, 5] are widely used to model sequential data in many areas such that natural language processing, speech recognition, machine translation, and stock market analysis. Given input and output data sets $\{\mathbf{a}_{i}\in\mathbb{R}^{\hat{n}}\}_{i=1}^{N}$ and $\{\mathbf{b}_{i}\in\mathbb{R}^{k}\}_{i=1}^{N}$, the training process of RNN is represented as
$\min_{\mathbf{x}}\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{b}_{i}-\mathbf{U}\sigma(\mathbf{W}(\cdots\sigma(\mathbf{V}\mathbf{a}_{1}+\mathbf{u})\cdots)+\mathbf{V}\mathbf{a}_{i}+\mathbf{u})+\mathbf{v})\|^{2},$ (2)
where $\mathbf{U}\in\mathbb{R}^{k\times l}$, ${\bf W}\in\mathbb{R}^{l\times l}$, ${\bf V}\in\mathbb{R}^{l\times\hat{n}}$ are the weight matrices, $\mathbf{u}\in\mathbb{R}^{l},\mathbf{v}\in\mathbb{R}^{k}$ are the bias vectors, $\sigma$ is an activation function, e.g., the ReLU activation function $\sigma({\bf z})=\max(0,{\bf z})$, and
$\mathbf{x}=({\rm vec}(\mathbf{U})^{\top},{\rm vec}(\mathbf{V})^{\top},{\rm vec}(\mathbf{W})^{\top},\mathbf{u}^{\top},\mathbf{v}^{\top})^{\top}\in\mathbb{R}^{n}$
with $n={kl+l^{2}+l\hat{n}+l+k}$ and $m=Nk$. When we study a big set of data, $m=Nk$ can be much bigger than $n$.
Problem (2) is an overdetermined least squares problem. Another example of overdetermined least squares problems is training convolutional neural network(CNN). For example, a two-layer CNN [3, 13]
$\min_{\mathbf{x}}\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{b}_{i}-\mathbf{U}\sigma(\mathbf{W}\sigma(\mathbf{V}\mathbf{a}_{i}+\mathbf{u})+\mathbf{v})\|^{2}.$ (3)
When the output data $\{\mathbf{b}_{i}\in\mathbb{R}^{k}\}_{i=1}^{N}$ contain noise, the following min-max problem
$\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y}):=\|\mathbf{F}(\mathbf{x})-\mathbf{C}\mathbf{y}\|^{2},$ (4)
is used for robust model training [17], where $\mathbf{C}\in\mathbb{R}^{m\times r}$ with $r\leq m$, $\Omega=\{\mathbf{y}\,:\,\|\mathbf{y}\|_{\infty}\leq\delta\}$, and $\delta\geq 0$. If $\delta=0$, then problem (4) reduces to problem (1).
First order methods including decent-ascent methods have been widely used to find the first order stationary points of min-max problems [4, 11, 10, 17]. However, the first order stationary points defined by gradients of the objective function $f$ can lead to incorrect solutions of min-max problems. In subsection 1.2, we show with an example of robust linear least square problems that solving the first order optimality conditions can lead to incorrect results.
This paper makes three contributions.
  • We use a change of variable in ${\bf y}$ to give an explicit formula for the value function of the inner maximization problem
    $\varphi(\mathbf{x})=\|\mathbf{F}(\mathbf{x})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}(\mathbf{x})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}=\max_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y}),$ (5)
    for any $\mathbf{x}\in\mathbb{R}^{n}$, where
    $\|{\bf C}\|_{F}=\left(\sum_{i=1,j=1}^{m,r}{\bf C}_{ij}^{2}\right)^{1/2}$
    is the Frobenius norm of ${\bf C}$.
  • We apply the formula for $\varphi$ to show that the method proposed by Gratton, Simon and Toint [9] can be used to find a first order $\epsilon$-approximate necessary minimax point of problem (4) with optimal complexity.
  • We give error bounds from any solution of nonlinear least-squares problem (1) to the set of minimizers of $\varphi.$
1.1 Important Concepts
In this subsection we review some subtle concepts for min-max problems. There are several forms of stationarity and they differ in important ways. These ideas will reappear throughout the paper.
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a saddle point [16] of problem (4) if for any $\mathbf{x}\in\mathbb{R}^{n},\mathbf{y}\in\Omega,$ we have
$f(\mathbf{x}^{*},\mathbf{y})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})\leq f(\mathbf{x},\mathbf{y}^{*}).$ (6)
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a local saddle point of problem (4) if (6) holds in a neighborhood of $(\mathbf{x}^{*},\mathbf{y}^{*})$.
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a minimax point [11] of problem (4) if
$\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})\leq\min_{\mathbf{x}}\max_{y\in\Omega}f(\mathbf{x},\mathbf{y}).$
A point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ is called a local minimax point of problem (4), if there exist a $\rho_{0}>0$ and a function $\tau:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ satisfying $\tau(\rho)\rightarrow 0$ as $\rho\rightarrow 0$, such that for any $\rho\in(0,\rho_{0}]$ and any $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{n}\times\Omega$ satisfying $\|\mathbf{x}-\mathbf{x}^{*}\|\leq\rho$ and $\|\mathbf{y}-\mathbf{y}^{*}\|\leq\rho$, we have
$f(\mathbf{x}^{*},\mathbf{y})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})\leq\max_{\mathbf{y}^{\prime}\in\{\mathbf{y}\in\Omega:\|\mathbf{y}-\mathbf{y}^{*}\|\leq\tau(\rho)\}}f(\mathbf{x},\mathbf{y}^{\prime}).$
By Theorem 3.11 in [10], if $f$ is directionally differentiable then a local minimax point $(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{R}^{n}\times\Omega$ of problem (4) satisfies the following first order condition
$\begin{array}[]{ll}d_{\mathbf{x}}f(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{v})=0,\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}\\ d_{\mathbf{y}}f(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{u})\leq 0,\quad\forall\,\mathbf{u}\in\mathrm{T}_{\Omega}(\mathbf{y}^{*}),\end{array}$ (7)
where $d_{\mathbf{x}}(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{v})$ and $d_{\mathbf{y}}(\mathbf{x}^{*},\mathbf{y}^{*};\mathbf{u})$ denote the directional derivatives of $f$ at $(\mathbf{x}^{*},\mathbf{y}^{*})$ along the directions $\mathbf{v}$ and $\mathbf{u}$, respectively, and $\mathrm{T}_{\Omega}(\mathbf{y})=:\left\{\mathbf{w}:\exists~{}\mathbf{y}^{k}\overset{\Omega}{\to}\mathbf{y},t^{k}\downarrow 0~{}\text{such that}~{}\lim_{k\to\infty}\frac{\mathbf{y}^{k}-\mathbf{y}}{t^{k}}=\mathbf{w}\right\}$ is the tangent cone to $\Omega$ at $\mathbf{y}\in\Omega$ [15, Definition 6.1]. When $\mathbf{F}$ is differentiable, the first order optimality condition of problem (4) has the following form
$\begin{array}[]{ll}\mathbf{F}^{\prime}(\mathbf{x})^{\top}(\mathbf{F}(\mathbf{x})-\mathbf{C}\mathbf{y})=0\\ \mathbf{y}={\rm Proj}_{\Omega}(\mathbf{y}-\alpha\mathbf{C}^{\top}(\mathbf{F}(\mathbf{x})-\mathbf{C}\mathbf{y})),\end{array}$ (10)
where $\alpha$ is any positive scalar.
Since the objective function is continuous and the feasible set $\Omega$ is bounded in problem (4), the solution set
$\Omega^{*}(\mathbf{x})=\mathop{\mathrm{argmax}}_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y})$
is nonempty and bounded for any $\mathbf{x}\in\mathbb{R}^{n}$. Hence, the function $\varphi$ is well-defined. However, it is known that solving a convex quadratic maximization over a box feasible set is NP-hard [14]. People often consider the first order optimality condition for nonconvex-nonconcave min-max problem [4, 10, 17]. However a solution of this first order optimality condition can be meaningless for (4).
1.2 Example
Consider problem (4) with $\mathbf{F}(\mathbf{x})=\mathbf{A}\mathbf{x}-\mathbf{b}$, $\mathbf{C}=\mathbf{I}$ and $n=m=r$. Suppose that $\mathbf{A}\in\mathbb{R}^{n\times n}$ is nonsingular. Then any pair $(\mathbf{x},\mathbf{y})=(\mathbf{A}^{-1}(\mathbf{b}+\mathbf{y}),\mathbf{y})$ with $\mathbf{y}\in\Omega$ satisfies the first order condition (10). On the other hand, the optimal solution of the inner maximization problem has the closed form
$\mathbf{y}_{i}(\mathbf{x})=\mathop{\mathrm{argmax}}_{|\mathbf{y}_{i}|\leq\delta}(\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y})_{i}^{2}=\delta\left\{\begin{array}[]{ll}1&{\rm if}\,(\mathbf{A}\mathbf{x}-\mathbf{b})_{i}<0\\ \{-1,1\}&{\rm if}\,(\mathbf{A}\mathbf{x}-\mathbf{b})_{i}=0\\ -1&{\rm if}\,(\mathbf{A}\mathbf{x}-\mathbf{b})_{i}>0,\end{array}\right.\quad i=1,\ldots,r.$ (11)
Hence
$\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}\|^{2}=\min_{\mathbf{x}}\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}(\mathbf{x})\|^{2}=r\delta^{2}$
at $(\mathbf{x}^{*},\hat{\mathbf{y}})=(\mathbf{A}^{-1}\mathbf{b},\delta\mathbf{c})$, where $\mathbf{c}$ is a vector with all elements being 1 or -1. In fact, by the Danskin theorem, the inner function $\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}(\mathbf{x})\|^{2}$ is strongly convex. Hence, the minimization problem has only one solution $\mathbf{x}^{*}=\mathbf{A}^{-1}\mathbf{b}.$ It is worth noting that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a saddle point of $f(\mathbf{x},\mathbf{y})=\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}\|^{2},$ for any $\hat{\mathbf{y}}$ with $\hat{\mathbf{y}}_{i}\in\{\delta,-\delta\}$, since
$f(\mathbf{x}^{*},\hat{\mathbf{y}})=r\delta^{2}>f(\mathbf{A}^{-1}(\mathbf{b}+\hat{\mathbf{y}}),\hat{\mathbf{y}})=0.$
In fact, $f$ does not have a saddle point, since
$r\delta^{2}=\min_{\mathbf{x}}\max_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y})\neq\max_{\mathbf{y}\in\Omega}\min_{\mathbf{x}}f(\mathbf{x},\mathbf{y})=0.$
See [6, Theorem 1.4.1]. Moreover, due to the arbitrary choice of $\delta$ for the closed form solution (11) of the maximization problem, we can easily find that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local saddle point.
On the other hand, we find
$\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y})=f(\mathbf{x}^{*},\hat{\mathbf{y}})=\min_{\mathbf{x}}\max_{y\in\Omega}f(\mathbf{x},\mathbf{y}).$
Hence $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is a minimax point of $f(\mathbf{x},\mathbf{y})=\|\mathbf{A}\mathbf{x}-\mathbf{b}-\mathbf{y}\|^{2}$. However, $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local minimax point of $f$. Indeed, consider $\hat{\mathbf{y}}=\delta\mathbf{e}$, where $\mathbf{e}$ is the vector with all elements being 1.
Let $\rho<\frac{\delta}{2\|{\bf A}\|}$ and $\tau(\rho)<\frac{\delta}{2}$ satisfying $\tau(\rho)\to 0$ as $\rho\to 0$. For any $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{n}\times\Omega$ satisfying $\|\mathbf{x}-\mathbf{x}^{*}\|\leq\rho$ and $\|\mathbf{y}-\delta\mathbf{e}\|\leq\rho$, we have
$|{\bf A}\mathbf{x}-\mathbf{b}|_{i}\leq\|{\bf A}\mathbf{x}-\mathbf{b}\|=\|{\bf A}(\mathbf{x}-\mathbf{x}^{*})\|\leq\|{\bf A}\|\rho\leq\delta-\frac{\tau(\rho)}{2}$
and for $i=1\ldots,r$,
[…]
Let $\mathbf{x}^{1}={\bf A}^{-1}(\mathbf{b}+\epsilon\mathbf{e})$ with $\epsilon\leq\rho/\|{\bf A}^{-1}\mathbf{e}\|$. Then we obtain $\|\mathbf{x}^{1}-\mathbf{x}^{*}\|\leq\rho$ and
$r\delta^{2}=f(\mathbf{x}^{*},\hat{\mathbf{y}})$ $>$ $\max_{\|\mathbf{y}\|_{\infty}\leq\delta,\|\mathbf{y}-\delta\mathbf{e}\|_{\infty}\leq\tau(\rho)}\|\mathbf{A}\mathbf{x}^{1}-\mathbf{b}-\mathbf{y}\|^{2}=\|\epsilon\mathbf{e}-\mathbf{y}(\mathbf{x})\|^{2}$
$=$ $(\epsilon-\delta)^{2}r\geq\min_{\mathbf{x}}\max_{\|\mathbf{y}\|_{\infty}\leq\delta,\|\mathbf{y}-\delta\mathbf{e}\|_{\infty}\leq\tau(\rho)}f(\mathbf{x},\mathbf{y}).$
Hence, $(\mathbf{x}^{*},\hat{\mathbf{y}})$ is not a local minimax point of $f$. Moreover, it is easy to see that $(\mathbf{x}^{*},\hat{\mathbf{y}})$ does not satisfy the first order condition (10), since ${\bf A}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b}-\hat{\mathbf{y}})=-\delta{\bf A}^{\top}\mathbf{e}\neq 0$. On the other hand, if $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ satisfies the first order condition (10), then from the nonsingularity of ${\bf A}$, we have $f(\mathbf{x}^{\prime},\mathbf{y}^{\prime})=0$, which implies $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ cannot be a local minimax point or a minimax point, since $00=f(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\not\geq\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{\prime},\mathbf{y})>$ 0.
Note that for nonconvex-nonconcave min-max optimization, a local minimax point must satisfy the first order condition, but a minimax point is not necessarily a stationary point[10].
This example shows that a min-max problem for robust linear least squares problems does not have a saddle point, a local saddle point and a local minimax point, while it has first order stationary points and minimax points. However, the set of first order stationary points and the set of minimax points do not have a common point.
This example also shows that the study of the value function of the inner maximization problem is important. In § 2, we give a proof of closed form expression (5) for the inner value function. Using the inner value function, we prove the existence of a minimax point of the robust nonlinear least squares problem (4). In § 3 we apply that formula to obtain a complexity bound for finding a first order stationary point for a minimax point of (4). Finally, in § 4 we apply the formula to expose some connections between solutions of (4) and those of (1), and give error bounds from any solution of nonlinear least squares problem (1) to the set of minimizers of $\varphi.$
2 The value function
The study of the inner function of the maximization problem is very important for solving min-max problem (4). Denote the value function by
$\varphi(\mathbf{x})=\max_{\mathbf{y}\in\Omega}f(\mathbf{x},\mathbf{y}).$
We consider a change of variables. Let
${\bf C}={\bf U}_{\bf C}\Sigma_{\bf C}{\bf V}^{\top}_{\bf C}$
be the singular value decomposition (SVD) of ${\bf C}$. Define
${\hat{\bf y}}={\bf V}^{\top}_{\bf C}{\bf y}\mbox{ and }{\hat{\bf C}}={\bf U}_{\bf C}\Sigma_{\bf C}.$
Then the QR factorization of ${\hat{\bf C}}$ is ${\bf U}_{\bf C}\Sigma_{\bf C}$ and the R factor is therefore diagonal. We argue that replacing ${\bf y}$ and ${\bf C}$ with ${\hat{\bf y}}$ and ${\hat{\bf C}}$ is consistent with the modeling assumption that the perturbation ${\bf C}{\bf y}$ be small and that rotating the coordinates for ${\bf y}$ does no harm. One can compare the new constraint $\|{\hat{\bf y}}\|_{\infty}\leq\delta$ with the constraint on ${\bf y}$ to see this by
$\|{\bf y}\|_{\infty}=\|{\bf V}_{\bf C}{\hat{\bf y}}\|_{\infty}\leq\|{\bf V}_{\bf C}\|_{\infty}\|{\hat{\bf y}}\|_{\infty}\leq\sqrt{r}\delta.$ (12)
So we will now assume that the change of variable has been made.
Assumption 2.1.
${\bf C}$ has full column rank. Let ${\bf C}={\bf Q}{\bf R}$ be the QR factorization of ${\bf C}$, where ${\bf R}$ is ($r\times r$) diagonal with positive diagonal entries.
We note that the orthogonality of ${\bf Q}$ implies that
$\|{\bf R}\|_{F}=\left(\sum_{j=1}^{r}{\bf R}_{jj}^{2}\right)^{1/2}=\|{\bf C}\|_{F}=\left(\sum_{i=1,j=1}^{m,r}{\bf C}^{2}_{ij}\right)^{1/2}.$
Hence (5) is equivalent to
$\varphi(\mathbf{x})=\|\mathbf{F}(\mathbf{x})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}(\mathbf{x})\|_{1}+\|{\bf R}\|_{F}^{2}\delta^{2}.$ (13)
We will prove (13). Our task is to compute the inner function
$\varphi(\mathbf{x})=\max_{{\bf y}\in\Omega}\|{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}.$
The orthogonal projection onto the range of ${\bf C}$ is
${\bf P}={{\bf Q}}{{\bf Q}}^{\top}$
which is an $m\times m$ matrix. If we project, since ${\bf P}$ is an orthogonal projector and ${\bf P}{\bf C}={\bf C}$ we have
$\|{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}=\|({\bf I}-{\bf P}){\bf F}({\bf x})\|^{2}+\|{\bf P}{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}$ (14)
by the Pythagorean theorem. The term $\|({\bf I}-{\bf P}){\bf F}({\bf x})\|^{2}$ does not depend on ${\bf y}$ so we have reduced the problem to computing
$\varphi_{0}({\bf x})=\max_{{\bf y}\in\Omega}\|{\bf P}{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}.$ (15)
Using the QR factorization again and the orthogonality of ${{\bf Q}}$ we get
$\begin{array}[]{ll}\|{\bf P}{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}&=\|{{\bf Q}}{{\bf Q}}^{\top}{\bf F}({\bf x})-{{\bf Q}}{{\bf R}}{\bf y}\|^{2}\\ \\ &=\|{{\bf Q}}^{\top}{\bf F}({\bf x})-{{\bf R}}{\bf y}\|^{2}.\end{array}$
Setting $\xi={{\bf Q}}^{\top}{\bf F}({\bf x})$ gives us the diagonal form of the maximization problem
$\varphi_{0}({\bf x})=\max_{{\bf y}\in\Omega}\|\xi-{{\bf R}}{\bf y}\|^{2}.$ (16)
We can solve this problem explicitly. Since ${\bf R}$ is diagonal with positive diagonal elements $R_{ii}$, we obtain
$\mathbf{y}_{i}=\delta\left\{\begin{array}[]{ll}-1&\mbox{if $\xi_{i}>0$}\\ \{-1,1\}&\mbox{if $\xi_{i}=0$}\\ 1&\mbox{if $\xi_{i}<0$,}\end{array}\right.\quad i=1,\ldots,r.$
Hence, from $\xi={\bf Q}^{\top}{\bf F}({\bf x})$, we have
$\varphi_{0}({\bf x})=\sum_{j=1}^{r}(|[{\bf Q}^{\top}{\bf F}({\bf x})]_{j}|+{\bf R}_{jj}\delta)^{2}.$ (17)
Here $[\cdot]_{i}$ denotes the $i$th component of a vector if the vector is a complicated expression, such as ${\bf Q}^{\top}{\bf F}({\bf x})$.
Since ${\bf P}={\bf Q}{\bf Q}^{\top}$ we have
$\sum_{j=1}^{r}|[{\bf Q}^{\top}{\bf F}({\bf x})]_{j}|^{2}=\|{\bf Q}^{\top}{\bf F}({\bf x})\|^{2}=\|{\bf P}{\bf F}({\bf x})\|^{2}$
and so, using the orthogonality of ${\bf P}$,
$\begin{array}[]{ll}\varphi({\bf x})&=\|({\bf I}-{\bf P}){\bf F}({\bf x})\|^{2}+\sum_{j=1}^{r}(|[{\bf Q}^{\top}{\bf F}({\bf x})]_{j}|+{\bf R}_{jj}\delta)^{2}\\ \\ &=\|({\bf I}-{\bf P}){\bf F}({\bf x})\|^{2}+\|{\bf P}{\bf F}({\bf x})\|^{2}+2\delta\sum_{j=1}^{r}{\bf R}_{jj}|[{\bf Q}^{\top}{\bf F}({\bf x})]_{j}|+(\sum_{j=1}^{r}{\bf R}_{jj}^{2})\delta^{2}\\ \\ &=\|{\bf F}({\bf x})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}+\|{\bf R}\|_{F}^{2}\delta^{2}.\\ \end{array}$ (18)
The last equality is a consequence of the diagonality of ${\bf R}$ and the fact that ${\bf R}{\bf Q}^{\top}={\bf C}^{\top}$. The term $\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}$ is nonsmooth, even if ${\bf F}$ is differentiable.
We summarize the results.
Theorem 2.1.
Let Assumption 2.1 hold. Then the inner function $\varphi$ is given by (5).
From the expression (5), the function $\varphi$ has the following properties.
If $\mathbf{F}$ is linear, then $\varphi$ is convex.
If $\mathbf{F}$ is continuously differentiable, then $\varphi$ is locally Lipschitz continuous, directionally differentiable, and its directional derivatives satisfy
$\varphi(\mathbf{x};\mathbf{d})=\max_{\mathbf{y}\in\Omega^{*}(\mathbf{x})}\mathbf{d}^{\top}{\bf F}^{\prime}(\mathbf{x})^{\top}(\mathbf{F}(\mathbf{x})-\mathbf{y}),$
[…]
If the set of minimizers of $\|\mathbf{F}(\mathbf{x})\|^{2}$ is nonempty and bounded, then the set of minimizers of $\varphi$ is nonempty and bounded, which can be seen from the relation of the two level sets
${\cal L}_{\varphi}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\varphi(\mathbf{x})\leq\beta\}\subseteq{\cal L}_{\mathbf{F}}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{F}(\mathbf{x})\|^{2}\leq\beta\},$
where $\beta>0$ is a scalar.
If $\|\mathbf{F}(\mathbf{x})\|^{2}$ has a minimizer $\mathbf{x}^{*}$ and there is $\gamma>0$ such that for any $\mathbf{x}\not\in{\bf X}_{\gamma}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{x}-\mathbf{x}^{*}\|\leq\gamma\}$, $\|{\bf C}\mathbf{F}(\mathbf{x})\|_{1}>\|{\bf C}\mathbf{F}(\mathbf{x}^{*})\|_{1}$, then $\varphi$ has a minimizer in ${\bf X}_{\gamma}$, which can be seen from the continuity of $\varphi$ and boundedness of ${\bf X}_{\gamma}$ with
$\varphi(\mathbf{x})-\varphi(x^{*})$ $=$ $\|\mathbf{F}(\mathbf{x})\|^{2}-\|\mathbf{F}(\mathbf{x}^{*})\|^{2}+2\delta(\|{\bf C}\mathbf{F}(\mathbf{x})\|_{1}-\|{\bf C}\mathbf{F}(\mathbf{x}^{*})\|_{1})$
$\geq$ $2\delta(\|{\bf C}\mathbf{F}(\mathbf{x})\|_{1}-\|{\bf C}\mathbf{F}(\mathbf{x}^{*})\|_{1})>0,\quad\mathbf{x}\not\in{\bf X}_{\gamma}$
and
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\varphi(\mathbf{x})\leq\min_{\mathbf{x}\in{\bf X}_{\gamma}}\varphi(\mathbf{x})\leq\varphi(\mathbf{x}^{*}).$
Theorem 2.2.
Let Assumption 2.1 hold. If $\varphi$ has a minimizer, then problem (4) has a minimax point and $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a minimax point of (4) for any
$\mathbf{x}^{*}\in\mathop{\mathrm{argmin}}_{{\bf x}}\varphi(\mathbf{x})\quad{\rm and}\quad\mathbf{y}^{*}\in\mathop{\mathrm{argmax}}_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y}).$
Proof 2.3.
Let $\mathbf{x}^{*}$ be a minimizer of $\varphi(\mathbf{x})$. Since $f$ is continuous and $\Omega$ is bounded, there is $\mathbf{y}^{*}$ such that
$f(\mathbf{x}^{*},\mathbf{y}^{*})=\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y}).$ (19)
Now we show $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a minimax point of (4). From (19), we have $\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y})\leq f(\mathbf{x}^{*},\mathbf{y}^{*}).$ On the other hand, we have
$\min_{\bf x}\max_{{\bf y}\in\Omega}f({\bf x},{\bf y})=\min_{\bf x}\varphi({\bf x})=\varphi({\bf x}^{*})=\max_{{\bf y}\in\Omega}f({\bf x}^{*},{\bf y})=f({\bf x}^{*},{\bf y}^{*}).$
Hence $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a minimax point of (4).
From Theorem 2.2, we can develop algorithms for finding a minimax point $(\mathbf{x}^{*},\mathbf{y}^{*})$ of (4) by the following two steps:
Step 1
Find $\mathbf{x}^{*}\in\mathop{\mathrm{argmin}}_{\bf x}\varphi(\mathbf{x}).$
Step 2
Find $\mathbf{y}^{*}\in\mathop{\mathrm{argmax}}_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y}).$
3 Complexity for minimax problem (4)
In this section, we show that the methods from [9] can be applied to find a first order $\epsilon$-approximate necessary minimax point of problem (4) with an optimal $O(|\log\epsilon|+\epsilon^{-2})$ evaluation bound. We assume that ${\bf F}$ is Lipschitz continuously differentiable in this section.
We consider the following minimization problem
$\min_{\bf x}\psi({\bf x}):=\|{\bf F}({\bf x})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}.$ (20)
Let the linearization of $\psi$ at $\mathbf{x}$ be defined as follows
$\phi({\bf x},{\bf s})=\|{\bf F}({\bf x})\|^{2}+2({\bf F}^{\prime}({\bf x})^{\top}{\bf F}({\bf x}))^{\top}{\bf s}+2\delta\|{\bf C}^{\top}({\bf F}({\bf x})+{\bf F}^{\prime}({\bf x}){\bf s})\|_{1}.$
From [9], one can verify that if ${\bf x}$ is a minimizer of problem (20) then
$\min_{\|{\bf s}\|\leq 1}\phi({\bf x},{\bf s})=\phi({\bf x},0).$ (21)
Following the definition in [9], we say that ${\bf x}$ is a first order $\epsilon$-approximate necessary minimizer of problem (20) if
$\phi({\bf x},0)-\min_{\|{\bf s}\|\leq 1}\phi({\bf x},{\bf s})=\max_{\|{\bf s}\|\leq 1}\Delta\phi({\bf x},{\bf s})\leq\epsilon,$ (22)
with
$\Delta\phi({\bf x},{\bf s})$ $=$ $\phi({\bf x},0)-\phi({\bf x},{\bf s})$
$=$ $-2({\bf F}^{\prime}({\bf x})^{\top}{\bf F}({\bf x}))^{\top}{\bf s}+2\delta(\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}-\|{\bf C}^{\top}({\bf F}({\bf x})+{\bf F}^{\prime}({\bf x}){\bf s})\|_{1}).$
We say that ${\bf x}$ is a first order necessary minimizer of problem (20) if $\epsilon=0$ in (22).
In [9], Gratton et al proposed an adaptive regularization algorithm for composite nonsmooth nonconvex optimization problems. Problem (20) satisfies Assumption AS.1-AS.4 in [9] when $\mathbf{F}$ is continuously differentiable and the level set $\{\mathbf{x}\,:\,\psi(\mathbf{x})\leq\psi(\mathbf{x}_{0})\}$ is bounded, where $\mathbf{x}_{0}$ is the initial point of the algorithm. Hence, we can use the algorithm to find a first order $\epsilon$-approximate necessary minimizer ${\bf x}$ satisfying (22), which needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of ${\bf F}({\bf x})$ and ${\bf F}^{\prime}({\bf x})$.
Definition 3.1.
We say $(\hat{{\bf x}},\hat{{\bf y}})$ is a first order $\epsilon$-approximate necessary minimax point of the min-max problem (4) if
$\min_{\|{\bf s}\|\leq 1}\Delta\phi(\hat{{\bf x}},{\bf s})\leq\epsilon,$ (23)
$\hat{\mathbf{y}}\in\mathop{\mathrm{argmax}}_{{\bf y}\in\Omega}\|{\bf F}(\hat{{\bf x}})-{\bf C}^{\top}{\bf y}\|^{2}.$ (24)
We say $(\hat{{\bf x}},\hat{{\bf y}})$ is a first order necessary minimax point of the minimax problem (4) if $\epsilon=0$ in (23).
Using the results in § 2, we can solve the maximization problem (24) explicitly. Let ${\bf C}={\bf Q}{\bf R}$. Then
$\hat{\mathbf{y}}_{i}=\delta\left\{\begin{array}[]{ll}-1&\mbox{if $({\bf Q}^{\top}{\bf F}(\hat{\mathbf{x}}))_{i}>0$}\\ \{-1,1\}&\mbox{if $({\bf Q}^{\top}{\bf F}(\hat{\mathbf{x}}))_{i}=0$}\\ 1&\mbox{if $({\bf Q}^{\top}{\bf F}(\hat{\mathbf{x}}))_{i}<0$,}\end{array}\right.\quad i=1,\ldots,r.$
Hence, by Theorem 2.2 and the complexity results for minimization problem [9], we obtain that finding a first order $\epsilon$-approximate necessary minimax point of (4) needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of ${\bf F}({\bf x})$ and ${\bf F}^{\prime}({\bf x})$.
4 Consequences of (5)
In this section, we use (5) to establish some interesting relations between the nonlinear least squares problem (1) and the robust nonlinear least squares problem (4). Denote the solution set of (1) and the set of minimizers of $\varphi$ defined in (5), respectively, by
${\bf X}^{*}:=\mathop{\mathrm{argmin}}_{\bf x}\,\|{\bf F}({\bf x})\|^{2}\quad{\rm and}\quad\hat{{\bf X}}:=\mathop{\mathrm{argmin}}_{\bf x}\,\varphi({\bf x}).$
We call ${\bf X}^{*}$ and $\hat{{\bf X}}$ are the nonlinear least squares solution set and the robust nonlinear least squares solution set, respectively. We call $\mathbf{x}^{*}\in{\bf X}^{*}$ and $\hat{\mathbf{x}}\in\hat{{\bf X}}$ are a nonlinear least squares solution and a robust nonlinear least squares solution, respectively.
4.1 $L^{1}$ optimality
Note that $\varphi({\hat{\bf x}})\leq\varphi({\bf x}^{*})$ and $\|{\bf F}({\bf x}^{*})\|^{2}\leq\|{\bf F}({\hat{\bf x}})\|^{2}$ for $\hat{\mathbf{x}}\in\hat{{\bf X}}$ and $\mathbf{x}^{*}\in{\bf X}^{*}$. Hence,
$\begin{array}[]{ll}\varphi({\hat{\bf x}})&=\|{\bf F}({\hat{\bf x}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({\hat{\bf x}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}\\ &\leq\varphi({{\bf x}^{*}})=\|{\bf F}({{\bf x}^{*}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}\\ &\leq\|{\bf F}({\hat{\bf x}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}.\end{array}$ (25)
Therefore
$\|{\bf C}^{\top}{\bf F}({\hat{\bf x}})\|_{1}\leq\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}.$ (26)
When ${\bf C}={\bf I}$, we can see that a least squares solution of (1) is not necessarily a least 1-norm solution of $\min_{\bf x}\|{\bf F}({\bf x})\|_{1}$.
When the equality holds in (26), then we have $\|{\bf F}({\hat{\bf x}})\|^{2}=\|{\bf F}({{\bf x}^{*}})\|^{2}$ and $\varphi({\hat{\bf x}})=\varphi({{\bf x}^{*}})$. Hence the nonlinear least squares solution set and the robust nonlinear least squares solution set are the same, that is ${\bf X}^{*}=\hat{{\bf X}}$. A special case is the zero residual problem.
If ${\bf F}({\bf x}^{*})=0$ then the nonsmooth term
$02\delta\|{\bf C}^{\top}{\bf F}({\bf x}^{*})\|_{1}=$ 0 .
Hence the two solution sets ${\bf X}^{*}$ and $\hat{{\bf X}}$ are the same. From
$\mathbf{y}^{*}\in\mathop{\mathrm{argmin}}_{{\bf y}\in\Omega}\|{\bf F}({\bf x}^{*})-{\bf C}^{\top}{\bf y}\|^{2}=\mathop{\mathrm{argmin}}_{{\bf y}\in\Omega}\|{\bf C}^{\top}{\bf y}\|^{2},$
we have ${\bf y}^{*}=\delta\mathbf{c}$, where $\mathbf{c}$ is a vector with all elements being 1 or -1 under Assumption 2.1. Summarizing, we have the following theorem.
Theorem 4.1.
Let Assumption 2.1 hold. If ${\bf F}({\bf x}^{*})=0$, then $({\bf x}^{*},\delta\mathbf{c})$ is a minimax point of (4), where $\mathbf{c}$ is a vector with all elements being 1 or -1.
However, as we saw in the example in subsection 1.1, $({\bf x}^{*},\delta\mathbf{c})$ need not be a saddle point, a local saddle point, a local minimax point or a stationary point of (4). Moreover, solutions of the first order conditions (7) or (10) may be not a minimax point of (4).
4.2 Error bounds: linear least squares problems
We first consider linear least squares problem,
${\bf F}({\bf x})={\bf A}{\bf x}-{\bf b}$
where ${\bf A}\in\mathbb{R}^{m\times n}$ and ${\bf b}\in\mathbb{R}^{m}$. Let
$\varphi_{\rm min}:=\min_{\bf x}\varphi(\mathbf{x}):=\|{\bf A}{\bf x}-{\bf b}\|^{2}+2\delta\|{\bf C}^{\top}({\bf A}{\bf x}-{\bf b})\|_{1}+\|{\bf C}\|^{2}_{F}\delta^{2}.$ (27)
Since the objective function of problem (27) is piecewise quadratic convex with nonnegative function values, problem (27) has a minimizer [7].
The subgradient of $\varphi$ at $\mathbf{x}$ is as follows
$\partial\varphi(\mathbf{x})=2{\bf A}^{\top}({\bf A}{\bf x}-{\bf b})+2\delta\partial\|{\bf C}^{\top}({\bf A}{\bf x}-{\bf b})\|_{1}.$
Let ${\bf B}={\bf C}^{\top}{\bf A}$ and ${\bf c}={\bf C}^{\top}{\bf b}$. Then
$\partial\|{\bf C}^{\top}({\bf A}{\bf x}-{\bf b})\|_{1}$ $=$ $\sum_{{\bf B}_{i}{\bf x}-{\bf c}_{i}>0}{\bf B}_{i}^{\top}-\sum_{{\bf B}_{i}{\bf x}-{\bf c}_{i}<0}{\bf B}_{i}^{\top}+\sum_{{\bf B}_{i}{\bf x}-{\bf c}_{i}=0}[\min({\bf B}_{i}^{\top},-{\bf B}_{i}^{\top}),\max({\bf B}_{i}^{\top},-{\bf B}_{i}^{\top})]$
$\subseteq$ $[-\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e},\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e}].$
where ${\bf B}_{i}$ is the $i$th row of ${\bf B}$.
Theorem 4.2.
For any given $\Gamma>0$, there is $\gamma>0$ such that for any ${\bf x}^{*}\in{\bf X}^{*}$ with $\varphi(\mathbf{x}^{*})-\varphi_{\rm min}\leq\Gamma$,
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq 2\delta\gamma^{2}\sqrt{r}\|{\bf C}^{\top}{\bf A}\|_{\infty}.$ (28)
Proof 4.3.
From Theorem 2.7 in [12], we have
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\gamma\sqrt{\varphi(\mathbf{x}^{*})-\varphi(\bar{{\bf x}})},$ (29)
if $\varphi(\mathbf{x}^{*})-\varphi_{\rm min}\leq\Gamma$.
Let $\bar{\mathbf{x}}\in\hat{{\bf X}}$ such that $\|\mathbf{x}^{*}-\bar{\mathbf{x}}\|={\rm dist}(\mathbf{x}^{*},\hat{{\bf X}}).$ Since $\varphi$ is convex, we have
$\varphi(\mathbf{x}^{*})-\varphi(\bar{\mathbf{x}})\leq g(\mathbf{x}^{*})^{\top}(\mathbf{x}^{*}-\bar{\mathbf{x}})\leq\|g(\mathbf{x}^{*})\|\|\mathbf{x}-\bar{\mathbf{x}}\|=\|g(\mathbf{x}^{*})\|{\rm dist}(\mathbf{x}^{*},\hat{{\bf X}}),$
where $g(\mathbf{x}^{*})\in\partial\varphi(\mathbf{x}^{*})$.
Since $\mathbf{x}^{*}\in{\bf X}^{*},$ we have $02{\bf A}^{\top}({\bf A}{\bf x}^{*}-{\bf b})=$ 0 and
$\partial\varphi(\mathbf{x}^{*})$ $=$ $2{\bf A}^{\top}({\bf A}{\bf x}^{*}-{\bf b})+2\delta\partial\|{\bf C}^{\top}({\bf A}{{\bf x}^{*}}-{\bf b})\|_{1}$
$=$ $2\delta\partial\|{\bf C}^{\top}({\bf A}{{\bf x}^{*}}-{\bf b})\|_{1}$
$\subseteq$ $2\delta[-\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e},\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e}].$
Hence from (29), we find
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\gamma^{2}\|g(\mathbf{x}^{*})\|\leq 2\delta\gamma^{2}\sqrt{r}\|{\bf C}^{\top}{\bf A}\|_{\infty}.$
This completes the proof.
From the first order necessary optimality conditions of (1) and (4) with ${\bf F}(\mathbf{x})={\bf A}\mathbf{x}-{\bf b}$, we have
${\bf A}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b})=0,\quad{\bf A}^{\top}({\bf A}\hat{\mathbf{x}}-{\bf b}-{\bf C}\mathbf{y})=0,$
for $\mathbf{x}^{*}\in{\bf X}^{*}$ and $\hat{\mathbf{x}}\in\hat{{\bf X}}$, which implies
${\bf A}^{\top}{\bf A}(\mathbf{x}^{*}-\hat{\mathbf{x}})={\bf C}\mathbf{y}.$
If ${\bf A}$ has full column rank, then
$\|\mathbf{x}^{*}-\hat{\mathbf{x}}\|\leq\|({\bf A}^{\top}{\bf A})^{-1}{\bf C}\|\|\mathbf{y}\|\leq\sqrt{r}\delta\|({\bf A}^{\top}{\bf A})^{-1}{\bf C}\|.$
Note that this error bound is the same $O(\delta)$ as (28). However, Theorem 4.2 does not assume that ${\bf A}$ has full column rank.
In the case of overdetermined linear least squares problems, with $m\gg n>r$, we often have $({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b}))_{i}\neq 0,$ for $i=1,\ldots,r$, where $\mathbf{x}^{*}$ is the minimum norm solution of the linear least squares problem
$\min\|{\bf A}\mathbf{x}-{\bf b}\|^{2}.$
Let the SVD of ${\bf A}$ be
${\bf A}={\bf U}_{A}\Sigma_{A}{\bf V}_{A}^{\top}.$
Then the pseudo-inverse of ${\bf A}$ is
${\bf A}^{\dagger}={\bf V}_{A}\Sigma_{A}^{\dagger}{\bf U}_{A}^{\top}$
and the minimum norm solution of the linear least squares problem is
${\bf x}^{*}={\bf A}^{\dagger}{\bf b}={\bf V}_{A}\Sigma_{A}^{\dagger}{\bf U}_{A}^{\top}{\bf b}=({\bf A}^{\top}{\bf A})^{\dagger}{\bf A}^{\top}{\bf b}.$
Let
$\eta=\min_{1\leq i\leq r}|{\bf C}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b})|_{i}.$ (30)
Proposition 4.4.
If $\eta>0$, then for any positive scalar $\delta$ satisfying
$\delta<\frac{\eta}{r\|{\bf C}^{\top}({\bf A}^{\top})^{\dagger}({\bf C}^{\top}{\bf A})^{\top}\|_{1}}$ (31)
the minimum norm solution of the robust linear least squares problem (27) is
$\hat{\mathbf{x}}=({\bf A}^{\top}{\bf A})^{\dagger}({\bf A}^{\top}{\bf b}-\delta\sum_{i=1}^{r}({\bf C}^{\top}{\bf A})^{\top}_{i}{\rm sgn}({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-b))_{i}),$ (32)
where $({\bf C}^{\top}{\bf A})_{i}$ is the $i$th row of ${\bf C}^{\top}{\bf A}$ and sgn is the sign function.
Proof 4.5.
First we show that $\min_{1\leq i\leq r}|{\bf C}^{\top}({\bf A}\hat{\mathbf{x}}-{\bf b})|_{i}\neq 0$, which implies that $\varphi$ is differentiable at $\hat{\mathbf{x}}$. From the SVD of ${\bf A}$ and the form of $\mathbf{x}^{*}$, we have
$\hat{\mathbf{x}}=\mathbf{x}^{*}-\delta({\bf A}^{\top}{\bf A})^{\dagger}\sum_{i=1}^{r}({\bf C}^{\top}{\bf A})^{\top}_{i}{\rm sgn}({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-b))_{i}.$ (33)
Hence for $i=1,\ldots,r$, we have
$|{\bf C}^{\top}({\bf A}\hat{\mathbf{x}}-{\bf b})-{\bf C}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b})|_{i}$ $=$ $\delta|{\bf C}^{\top}{\bf A}({\bf A}^{\top}{\bf A})^{\dagger}\sum_{i=1}^{r}({\bf C}^{\top}{\bf A})^{\top}_{i}{\rm sgn}({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-b))_{i}|_{i}$
$=$ $\delta|{\bf C}^{\top}({\bf A}^{\top})^{\dagger}\sum_{i=1}^{r}({\bf C}^{\top}{\bf A})^{\top}_{i}{\rm sgn}({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-b))_{i}|_{i}$
$\leq$ $\delta\sum_{i=1}^{r}|{\bf C}^{\top}({\bf A}^{\top})^{\dagger}({\bf C}^{\top}{\bf A})^{\top}_{i}|_{i}$
$\leq$ $\delta r\|{\bf C}^{\top}({\bf A}^{\top})^{\dagger}({\bf C}^{\top}{\bf A})^{\top}\|_{1}$
$<$ $\eta.$
By the definition of $\eta$ in (30), we find that for $i=1,\ldots,r$, $|{\bf C}^{\top}({\bf A}\hat{\mathbf{x}}-{\bf b})|_{i}\neq 0$ and
${\rm sin}({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-b))_{i}={\rm sin}({\bf C}^{\top}({\bf A}\hat{\mathbf{x}}-b))_{i}.$
Now we show $\hat{\mathbf{x}}$ is a minimum norm minimizer of $\varphi$. Since $\varphi$ is a convex function and differentiable at $\hat{\mathbf{x}}$, the first order optimality condition is necessary and sufficient for the optimality, which holds at $\hat{\mathbf{x}}$ as the follows
$\nabla\varphi(\hat{\mathbf{x}})$ $=$ $2{\bf A}^{\top}({\bf A}\hat{\mathbf{x}}-b)+2\delta\sum_{i=1}^{r}({\bf C}^{\top}{\bf A})_{i}{\rm sgn}({\bf C}^{\top}({\bf A}\hat{\mathbf{x}}-b))_{i}$
$=$ $2{\bf A}^{\top}{\bf A}\hat{\mathbf{x}}-2{\bf A}^{\top}b+2\delta\sum_{i=1}^{r}({\bf C}^{\top}{\bf A})_{i}{\rm sgn}({\bf C}^{\top}({\bf A}\mathbf{x}^{*}-b))_{i}$
$=$ $0,$
where we use (32). Hence $\hat{\mathbf{x}}$ is the minimum norm solution of (27).
Remark 4.6.
If $\mathbf{F}(\mathbf{x}^{*})={\bf A}\mathbf{x}^{*}-{\bf b}\neq 0$, then we can choose a matrix ${\bf C}\in\mathbb{R}^{m\times r}$ with rank$({\bf C})=r$ such that $\eta=\min_{1\leq i\leq r}|{\bf C}^{\top}({\bf A}\mathbf{x}^{*}-{\bf b})|_{i}>0$. For example, let
${\bf C}_{i,j}=\left\{\begin{array}[]{lr}1&{\rm if}\,\,\mathbf{F}_{i}(\mathbf{x}^{*})\geq 0,\,\,i\neq j\\ -1&{\rm if}\,\,\mathbf{F}_{i}(\mathbf{x}^{*})<0,\,\,i\neq j\\ 2r&{\rm if}\,\,\mathbf{F}_{i}(\mathbf{x}^{*})\geq 0,\,\,i=j\\ -2r&{\rm if}\,\,\mathbf{F}_{i}(\mathbf{x}^{*})<0,\,\,i=j\end{array}\right.$
for $i=1,\ldots,m$ and $j=1,\ldots,r$. Then $\eta\geq\|\mathbf{F}(\mathbf{x}^{*})\|_{1}$. Moreover, the submatrix of the first $r$ rows of ${\bf C}$ is strictly diagonally dominant. Hence rank$({\bf C})=r$.
From (33), we also obtain an error bound
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\|\mathbf{x}^{*}-\hat{\mathbf{x}}\|\leq 2\delta r\|({\bf A}^{\top}{\bf A})^{\dagger}({\bf C}^{\top}{\bf A})^{\top}\|_{1}$
in a neighborhood of the minimum norm solution $\mathbf{x}^{*}$ of the linear least squares problem when $\varphi$ is differentiable at $\mathbf{x}^{*}$. This special case is different from the one in Equation (24) (page 1044) of [8]. Note that in [8], the set $\Omega=\{\mathbf{y}\,:\,\|\mathbf{y}\|\leq\delta\}$ is defined by the Euclidean norm and a linear least squares solution is a robust linear least squares solution.
4.3 Error bounds: nonlinear least squares problems
We use the second order growth condition in [1, Section 4.4.1] to present an error bound from a nonlinear least squares solution to the robust nonlinear least squares solution set.
We say that the second order growth condition holds at $\hat{{\bf X}}$ if there exists a neighborhood ${\cal N}$ of $\hat{{\bf X}}$ and a constant $\mu>0$ such that
$\varphi(\mathbf{x})\geq\varphi_{\rm min}+\mu[{\rm dist}(\mathbf{x},\hat{{\bf X}})]^{2},\,\,\,\forall\mathbf{x}\,\in{\cal N},$ (34)
where $\varphi_{\rm min}=\min_{\mathbf{x}}\varphi(\mathbf{x}).$ If $\varphi$ is a piecewise convex quadratic function, then (34) holds [12].
We say that ${\bf C}\mathbf{F}$ is Lipschitz continuous in ${\cal N}$ at $\hat{{\bf X}}$ if there is $L>0$ such that
$\|{\bf C}(\mathbf{F}(\mathbf{x})-\mathbf{F}(\mathbf{z}))\|\leq L\|\mathbf{x}-\mathbf{z}\|,\quad\quad\forall\mathbf{x}\,\in{\cal N},\mathbf{z}\in\hat{{\bf X}}.$ (35)
Theorem 4.7.
Suppose the second order growth condition (34) and the Lipschitz continuity condition (35) hold. Then for any $\mathbf{x}^{*}\in{\bf X}^{*}\cap{\cal N}$, we have
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq 2\delta\sqrt{r}L\mu^{-1}.$ (36)
Proof 4.8.
Let $\hat{\mathbf{x}}\in\hat{{\bf X}}$ such that $\|\mathbf{x}^{*}-\hat{\mathbf{x}}\|=$dist$(\mathbf{x}^{*},\hat{{\bf X}}).$ From condition (35), we have
$\varphi(\mathbf{x}^{*})-\varphi(\hat{\mathbf{x}})$ $=$ $\|\mathbf{F}(\mathbf{x}^{*})\|+2\delta\|{\bf C}^{\top}\mathbf{F}(\mathbf{x}^{*})\|_{1}-\|\mathbf{F}(\hat{\mathbf{x}})\|-2\delta\|{\bf C}^{\top}\mathbf{F}(\hat{\mathbf{x}})\|_{1}$
$\leq$ $2\delta\|{\bf C}^{\top}\mathbf{F}(\mathbf{x}^{*})\|_{1}-2\delta\|{\bf C}^{\top}\mathbf{F}(\hat{\mathbf{x}})\|_{1}$
$\leq$ $2\delta\|{\bf C}^{\top}(\mathbf{F}(\mathbf{x}^{*})-\mathbf{F}(\hat{\mathbf{x}}))\|_{1}$
$\leq$ $2\delta\sqrt{r}\|{\bf C}^{\top}(\mathbf{F}(\mathbf{x}^{*})-\mathbf{F}(\hat{\mathbf{x}}))\|$
$\leq$ $2\delta\sqrt{r}L\|\mathbf{x}^{*}-\hat{\mathbf{x}}\|$
$=$ $2\delta\sqrt{r}L{\rm dist}(\mathbf{x}^{*},\hat{{\bf X}}),$
where the first inequality uses $\|\mathbf{F}(\mathbf{x}^{*})\|\leq\|\mathbf{F}(\hat{\mathbf{x}})\|.$ Together with the second order growth condition (34), we find the error bound (36).
5 Conclusion
We consider the min-max problem (4) for robust nonlinear least squares problems. We give an explicit formula (5) for the value function of the inner maximization problem in (4). Using the formula (5), we show that finding a first order $\epsilon$-approximate necessary minimax point of the min-max problem needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of the function value and gradient of the objective function. Moreover, we provide error bounds from a solution of the least squares problem (1) to the solution set of the robust least squares problem.